OpenStudy (anonymous):

Solve: | m + 8| = 16 {-24, 8} {8, 16} {-24}

OpenStudy (ash2326):

Absolute value function is defined as |x| = x for x>=0 -x for x<0 here we have \[|m+8|=16\] so \[ \pm(m+8)=16\] or \[m+8=\pm 16\] \[ m+8=16\ or \ m+8=-16\] we get \[m=8\ or\ m=-24\] so m= {-24, 8}